OpenStudy (anonymous):

X to the 4rth power (times) x to the 6th power=

OpenStudy (anonymous):

when you forget how to solve it, make yourself an example 2^2 * 2^3 = 4 * 8 = 32, and 2^5=32, then you discover 2+3=5, so the power adds together same logic, x^4 * x^6 = x^(4+6) = x^10